OpenStudy (anonymous):

PLEASE HELP 1) 42 is what percent of 63? the answer is: 66 2/3% but i don't know the steps so can someone explain the steps to me?

OpenStudy (mathstudent55):

A percent is a comparison of numbers using a ratio whose denominator is 100. For example, one-half can be written as 1/2. If you want to know what 1/2 is as a percent, make the denominator of 1/2 become 100. Then the numerator will be the percent. \(\dfrac{1}{2} = \dfrac{1}{2} \times \dfrac{50}{50} = \dfrac{50}{100} \) Notice that the last fraction, which equals 1/2 is 50/100. Since now you have a ratio based on 100, then the percent is 50%. That means 1/2 is the same as 50%

OpenStudy (mathstudent55):

What is 3/4 as a percent? Find an equivalent fraction to 3/4 with 100 as the denominator. \(\dfrac{3}{4} = \dfrac{3}{4} \times \dfrac{25}{25} = \dfrac{75}{100} \) Now you have 75/100 which is an equivalent fraction to 3/4, and since 75/100 has a denominator of 100, so 3/4 is the same as 75%.

OpenStudy (mathstudent55):

Now we can do your problem. You want to know what percent 42 is of 63. First, let's just write it as a ratio: \(\dfrac{42}{63} \) We need the denominator to be 100 to have a percentage. \(\dfrac{42}{63} = \dfrac{x}{100} \) We are asking the question: 42 out of 63 is how much out of 100? Something out of 100 is a percent. We can solve the proportion above to find x, the percent. \(\dfrac{42}{63} = \dfrac{x}{100}\) We cross multiply: \(63x = 42 \times 100\) \(63x = 4200\) Divide both sides by 63: \(\dfrac{63x}{63} = \dfrac{4200}{63} \) Using a calculator to do 6300/42, we get: x = 66.666666... 66.6666... = 66 2/3 That means that 42 out of 63 is 66 2/3%